OpenStudy (anonymous):

Solve for x 1/3 (2x - 8) = 4. (1 point) 2 6 10 10/3

OpenStudy (texaschic101):

1/3(2x - 8) = 4 --- multiply both sides by 3 (the reciprocal) 2x - 8 = 12 -- add 8 to both sides 2x = 12 + 8 2x = 20 x = 10